# Molecular Quantum Computations on a Protein

### v1

@@ 3.1.2 Mean-field calculations
All mean-field calculations for the folded and unfolded Trp-cage conformers were performed using GPU4PySCF li2024introducting, wu2024enhancing employing the restricted HF method with the STO-3G basis set (HF/STO-3G). Density fitting B000027M was enabled throughout to reduce the overhead of the integral evaluation and maximize throughput on GPU hardware. The default PySCF auxiliary basis applied during density fitting was def2-SVP-JKFIT for the elements present in Trp-cage (H, C, N, and O). Each calculation was executed on a single NVIDIA V100S-PCIe 32 GB accelerator paired with a 2.4 GHz Intel Xeon Platinum 8260 CPU. The GPU-accelerated Hartree-Fock evaluations finished in roughly four minutes per conformer. This mean-field stage establishes the orbital and density matrix input that feeds into the subsequent fragmentation, bath construction, and correlated solver workflows.
@@ 3.2.2 Quantum circuit design: the LUCJ ansatz
where $\hat{a}^{\dagger}$ ($\hat{a}$) are creation (annihilation) operators, $p$,$r$,$s$, and $q=1\dots M$ denote basis set elements (MOs), $\sigma$ and $\tau$ denote spin-$z$ polarizations, $h_{pr}$ and $(pr|qs)$ are the one- and two-body electronic integrals, and $E_{0}$ is a constant accounting for the electrostatic interactions between nuclei and electrons in occupied inactive orbitals. We obtain the quantities $E_{0}$, $h_{pr}$, and $(pr|qs)$ for the selected active spaces using PySCF.sun2020recent, sun2018pyscf, sun2015libcint
We prepare our quantum circuits, used to sample configurations, starting from a truncated version of the local unitary cluster Jastrow (LUCJ) ansatz motta2023bridging
@@
where ${\hat{K}_{\mu}}=\Sigma_{pr,\sigma}K^{\mu}_{pr}\hat{a}_{p\sigma}^{\dagger}\hat{a}^{\phantom{\dagger}}_{r\sigma}$ are one-body operators, ${\hat{J}_{\mu}}=\Sigma_{pr,\sigma\tau}J^{\mu}_{p\sigma,r\tau}\hat{n}_{p\sigma}\hat{n}_{r\tau}$ are density-density operators, and $|{\bf{x}}_{\mathrm{RHF}}\rangle$ is the restricted closed-shell Hartree-Fock (RHF) state. We use the Jordan-Wigner (JW) transformation ortiz2002simulating to map the fermionic wavefunction Eq. (2) onto a qubit wavefunction that can be prepared executing a quantum circuit. The JW transformation maps the Fock space of fermions in $M$ spatial orbitals onto the Hilbert space of $2M$ qubits, where the basis state $|{\bf{x}}\rangle$ is parametrized by a bitstring ${\bf{x}}\in\{0,1\}^{2M}$ and represents an electronic configuration where the spin-orbital $p\sigma$ is occupied (empty) if $x_{p\sigma}=1$ ($x_{p\sigma}=0$). We prepare the wavefunction given in Eq. (2) by executing the following quantum circuit: a single layer of Pauli-X gates prepares the basis state $|{\bf{x}}_{\mathrm{RHF}}\rangle$, a Bogoliubov circuit aleksandrowicz2019qiskit (with linear depth, quadratic number of gates, and a 1D qubit connectivity) encodes each orbital rotation $e^{\pm\hat{K}_{\mu}}$, and a circuit of Pauli-ZZ rotations encodes each density-density interaction $e^{i\hat{J}_{\mu}}$. When $J^{\mu}$ is a dense matrix, Pauli-ZZ rotations are applied across all pair of qubits, requiring all-to-all qubit connectivity or a substantial overhead of SWAP gates. To mitigate these quantum hardware requirements LUCJ imposes a “locality” approximation, e.g., in the most extreme form it enforces $J^{\mu}_{p\sigma,r\tau}=0$ for all pairs of spin-orbitals that are not mapped onto adjacent qubits under JW motta2023bridging (as a consequence, a circuit with constant depth and linear number of gates encodes each $e^{i\hat{J}_{\mu}}$ operator). Furthermore, to control the depth and size of quantum circuits, the number of layers in Eq. (2) is $L=1$. We produce the LUCJ circuits using the ffsim library ffsim2024 interfaced with Qiskit 1.1.1 aleksandrowicz2019qiskit, javadi2024quantum.
@@
LUCJ circuits can be initialized using CCSD t-amplitudes via a double-factorized representation motta2021low of the $\hat{T}_{2}$ operators. We rewrite $\hat{T}_{2}-\hat{T}_{2}^{\dagger}$ in unitary CCSD theory as
@@
Parameter Optimization: We optimize the LUCJ parameters according to the guidelines in Lin et al lin2025pushing. A recent study showed significant improvements of the sampling ability upon optimization of the LUCJ parameters Lin_2025_ImprovedParameterInitialization. The performance of the LUCJ circuits is improved by applying a compressed double factorization to the $t_{2}$-amplitudes of $\hat{T}_{2}$ operator:
@@
where $\bar{t}_{ij}^{ab}$ are as in Eq. (7) and $t_{ij}^{ab}$ are the original CCSD $t_{2}$-amplitudes. Optimizations employ L-BFGS-B in SciPy with gradients obtained using automatic differentiation. In the present study we use single-stage optimization due to the fact that multi-stage optimization Lin_2025_ImprovedParameterInitialization is computationally expensive for the largest EWF-cluster.
@@ 3.2.3 Sampling using circuits run on quantum hardware
For the hardware simulations, all circuits were executed on IBM Heron-R2 QPUs with 1,000,000 measurement outcomes (shot) per circuit. LUCJ circuits with $L=1$ were used and four $\alpha$$\beta$ qubits were connected via ancillary qubits (see Fig. 6 for the qubit layout). Among the available zig-zag connectivity patterns, the layout was selected using a heuristic that sums the two-qubit gate errors of all couplers and the readout errors of the qubits in the pattern IBM2025SQD.
Qiskit v2.1.1javadi2024quantum was used with generate_preset_pass_managergenerate\_preset\_pass\_manager at optimization_level=0optimization\_level=0, incorporating ffsim’s PRE_INITPRE\_INIT passes for orbital-rotation decomposition and applying $RemoveIdentityEquivalent$ post-initialization to remove negligible (i.e. near-identity) rotations. Dynamic decoupling using the $XY4$ sequence was applied to further suppress decoherence.
@@ 3.2.4 Ground State and Energy of the Hamiltonian in SQD
We use a technique called self-consistent configuration recovery robledo2024chemistry, executed on a classical computer, to restore particle number conservation. The associated code is publicly available in the GitHub repository. sqd_addon Within each step of self-consistent recovery, we sample $K$ subsets (or batches) of $\tilde{\chi}$ labeled $\tilde{\chi}_{b}$ with $b=1\dots K$. Each batch defines – through a transformation robledo2024chemistry informed by an approximation to the ground-state occupation numbers $n_{p\sigma}$ – a subspace $S^{(b)}$ of dimension $d$, in which we project the many-electron Hamiltonian as kanno2023quantum, nakagawa2023adapt, robledo2024chemistry
@@
Moreover, we set the convergence thresholds for total energy and average orbital occupancy as $1\cdot 10^{-8}$ and $1\cdot 10^{-5}$ respectively, and we set the maximum number of configuration recovery iterations at 5. We use “Qiskit Addon: SQD Version 0.12.0” in which we select the most relevant electron configurations, based on their configuration interaction coefficients at each step of the configuration recovery procedure. These selected configurations are always included in all the batches produced. Shirakawa_2025_ClosedLoopQuantumClassical The threshold for identifying configurations as relevant was set to $1\cdot 10^{-4}$. All simulations were performed with the spin symmetrization for $\alpha$ and $\beta$ spin-orbitals. In the first iteration of self-consistent configuration recovery, we initialize $n_{p\sigma}$ from measurement results in $\tilde{\chi}$ with the correct particle number.
The diagonalization in each step of configuration recovery was performed with the SBD solver which can be obtained through the GitHub repository.sbd The diagonalization in the SBD solver is parallelized with 48 CPUs used in the diagonalization of each subspace, where we use a total of 480 CPUs in 10 batches for each EWF cluster. The full sequence of configuration recovery, subspace construction, and diagonalization steps is summarized schematically in Fig. 8.
Subspace augmentation (ext-SQD): The extended SQD (ext-SQD) procedure was introduced to enable excited-state calculations based on SQD barison2025quantum, shivpuje2025sample but may also be used to improve the accuracy of ground- and excited-state SQD calculations barroca2025surface.
In ext-SQD we first take the electron configurations from the lowest energy batch on the last step of the configuration recovery procedure. We select the most dominant electron configurations based on the value of their configuration interaction coefficients (dominant if the threshold is above $1\cdot 10^{-5}$). Using the selected configurations, we then augment their subspaces by applying single excitations on each configuration within the selected set. The augmentation of the subspace was peformed with the PyCI software package.richer2024pyci We use the augmented subspace to perform diagonalization within the given EWF cluster to obtain the ground state and energy of the Hamiltonian. Finally, we use the ground state produced after the diagonalization of the augmented subspace to calculate the reduced density matrices (RDM) of each EWF cluster.
@@ 5 Conclusions and Outlook
This study presents two important breakthroughs. First, the simulation of Trp-cage with STO-3G basis set is the largest example of embedding-based simulations within the configuration interaction paradigm and to our knowledge this is the largest case ever simulated exclusively within the configuration interaction framework. Trp-cage simulated with the STO-3G basis set results in simulation of 919 molecular orbitals (MOs), where the largest DMET and EWF simulations were reported only for systems explicitly including only up to 321 and 312 MOs, respectively.wouters2016practical, nusspickel2022systematic Second, this represents the first demonstration of a system of this biochemical size being treated with quantum sampling on a quantum device as part of an embedding workflow, illustrating both the feasibility and the growing capability of hybrid quantum–classical electronic structure methods.
The present study paves the way for fragment-based solutions in quantum computing. Our approach is highly modular and allows for integration of novel quantum computing methodologies in the future. Hence, the present workflow will gradually mature in a synchronous manner with the maturation of quantum hardware and quantum algorithms. In the long-term, upon establishment of fault-tolerant quantum computing, the present fragment-based methodology can eventually adapt even more powerful fully-quantum methods such as quantum phase estimation (QPE) where perspectives of utilization of this method within embedding schemes were recently explored by Erakovic et al.erakovic2025high
Configuration interaction simulations accelerated with quantum computers can offer unique opportunities for simulations of complex chemical processes in proteins such as photochemical reactions and proton-coupled electron transfermai2020molecular, warburton2022theoretical. In the long-term, upon adaptation of more advanced future quantum hardware and quantum algorithms, this methodology potentially might offer a unified highly-accurate method for the treatment of any protein-based system beyond the scaling of classical computers.

### v2

@@ 3.1.2 Mean-field calculations
All mean-field calculations for the folded and unfolded Trp-cage conformers were performed using GPU4PySCF 38, 97 employing the restricted HF method with the STO-3G basis set (HF/STO-3G). Density fitting 16 was enabled throughout to reduce the overhead of the integral evaluation and maximize throughput on GPU hardware. The default PySCF auxiliary basis applied during density fitting was def2-SVP-JKFIT for the elements present in Trp-cage (H, C, N, and O). Each calculation was executed on a single NVIDIA V100S-PCIe 32 GB accelerator paired with a 2.4 GHz Intel Xeon Platinum 8260 CPU. The GPU-accelerated Hartree-Fock evaluations finished in roughly four minutes per conformer. This mean-field stage establishes the orbital and density matrix input that feeds into the subsequent fragmentation, bath construction, and correlated solver workflows.
@@ 3.2.2 Quantum circuit design: the LUCJ ansatz
where $\hat{a}^{\dagger}$ ($\hat{a}$) are creation (annihilation) operators, $p$,$r$,$s$, and $q=1\dots M$ denote basis set elements (MOs), $\sigma$ and $\tau$ denote spin-$z$ polarizations, $h_{pr}$ and $(pr|qs)$ are the one- and two-body electronic integrals, and $E_{0}$ is a constant accounting for the electrostatic interactions between nuclei and electrons in occupied inactive orbitals. We obtain the quantities $E_{0}$, $h_{pr}$, and $(pr|qs)$ for the selected active spaces using PySCF.87, 86, 88
We prepare our quantum circuits, used to sample configurations, starting from a truncated version of the local unitary cluster Jastrow (LUCJ) ansatz 49
@@
where ${\hat{K}_{\mu}}=\Sigma_{pr,\sigma}K^{\mu}_{pr}\hat{a}_{p\sigma}^{\dagger}\hat{a}^{\phantom{\dagger}}_{r\sigma}$ are one-body operators, ${\hat{J}_{\mu}}=\Sigma_{pr,\sigma\tau}J^{\mu}_{p\sigma,r\tau}\hat{n}_{p\sigma}\hat{n}_{r\tau}$ are density-density operators, and $|{\bf{x}}_{\mathrm{RHF}}\rangle$ is the restricted closed-shell Hartree-Fock (RHF) state. We use the Jordan-Wigner (JW) transformation 59 to map the fermionic wavefunction Eq. (2) onto a qubit wavefunction that can be prepared executing a quantum circuit. The JW transformation maps the Fock space of fermions in $M$ spatial orbitals onto the Hilbert space of $2M$ qubits, where the basis state $|{\bf{x}}\rangle$ is parametrized by a bitstring ${\bf{x}}\in\{0,1\}^{2M}$ and represents an electronic configuration where the spin-orbital $p\sigma$ is occupied (empty) if $x_{p\sigma}=1$ ($x_{p\sigma}=0$). We prepare the wavefunction given in Eq. (2) by executing the following quantum circuit: a single layer of Pauli-X gates prepares the basis state $|{\bf{x}}_{\mathrm{RHF}}\rangle$, a Bogoliubov circuit 1 (with linear depth, quadratic number of gates, and a 1D qubit connectivity) encodes each orbital rotation $e^{\pm\hat{K}_{\mu}}$, and a circuit of Pauli-ZZ rotations encodes each density-density interaction $e^{i\hat{J}_{\mu}}$. When $J^{\mu}$ is a dense matrix, Pauli-ZZ rotations are applied across all pair of qubits, requiring all-to-all qubit connectivity or a substantial overhead of SWAP gates. To mitigate these quantum hardware requirements LUCJ imposes a “locality” approximation, e.g., in the most extreme form it enforces $J^{\mu}_{p\sigma,r\tau}=0$ for all pairs of spin-orbitals that are not mapped onto adjacent qubits under JW 49 (as a consequence, a circuit with constant depth and linear number of gates encodes each $e^{i\hat{J}_{\mu}}$ operator). Furthermore, to control the depth and size of quantum circuits, the number of layers in Eq. (2) is $L=1$. We produce the LUCJ circuits using the ffsim library 21 interfaced with Qiskit 1.1.1 1, 29.
@@
LUCJ circuits can be initialized using CCSD t-amplitudes via a double-factorized representation 51 of the $\hat{T}_{2}$ operators. We rewrite $\hat{T}_{2}-\hat{T}_{2}^{\dagger}$ in unitary CCSD theory as
@@
Parameter Optimization: We optimize the LUCJ parameters according to the guidelines in Lin et al 43. A recent study showed significant improvements of the sampling ability upon optimization of the LUCJ parameters 42. The performance of the LUCJ circuits is improved by applying a compressed double factorization to the $t_{2}$-amplitudes of $\hat{T}_{2}$ operator:
@@
where $\bar{t}_{ij}^{ab}$ are as in Eq. (7) and $t_{ij}^{ab}$ are the original CCSD $t_{2}$-amplitudes. Optimizations employ L-BFGS-B in SciPy with gradients obtained using automatic differentiation. In the present study we use single-stage optimization due to the fact that multi-stage optimization 42 is computationally expensive for the largest EWF-cluster.
@@ 3.2.3 Sampling using circuits run on quantum hardware
For the hardware simulations, all circuits were executed on IBM Heron-R2 QPUs with 1,000,000 measurement outcomes (shot) per circuit. LUCJ circuits with $L=1$ were used and four $\alpha$$\beta$ qubits were connected via ancillary qubits (see Fig. 6 for the qubit layout). Among the available zig-zag connectivity patterns, the layout was selected using a heuristic that sums the two-qubit gate errors of all couplers and the readout errors of the qubits in the pattern 27.
Qiskit v2.1.129 was used with generate_preset_pass_managergenerate\_preset\_pass\_manager at optimization_level=0optimization\_level=0, incorporating ffsim’s PRE_INITPRE\_INIT passes for orbital-rotation decomposition and applying $RemoveIdentityEquivalent$ post-initialization to remove negligible (i.e. near-identity) rotations. Dynamic decoupling using the $XY4$ sequence was applied to further suppress decoherence.
@@ 3.2.4 Ground State and Energy of the Hamiltonian in SQD
We use a technique called self-consistent configuration recovery 70, executed on a classical computer, to restore particle number conservation. The associated code is publicly available in the GitHub repository. 73 Within each step of self-consistent recovery, we sample $K$ subsets (or batches) of $\tilde{\chi}$ labeled $\tilde{\chi}_{b}$ with $b=1\dots K$. Each batch defines – through a transformation 70 informed by an approximation to the ground-state occupation numbers $n_{p\sigma}$ – a subspace $S^{(b)}$ of dimension $d$, in which we project the many-electron Hamiltonian as 32, 53, 70
@@
Moreover, we set the convergence thresholds for total energy and average orbital occupancy as $1\cdot 10^{-8}$ and $1\cdot 10^{-5}$ respectively, and we set the maximum number of configuration recovery iterations at 5. We use “Qiskit Addon: SQD Version 0.12.0” in which we select the most relevant electron configurations, based on their configuration interaction coefficients at each step of the configuration recovery procedure. These selected configurations are always included in all the batches produced. 79 The threshold for identifying configurations as relevant was set to $1\cdot 10^{-4}$. All simulations were performed with the spin symmetrization for $\alpha$ and $\beta$ spin-orbitals. In the first iteration of self-consistent configuration recovery, we initialize $n_{p\sigma}$ from measurement results in $\tilde{\chi}$ with the correct particle number.
The diagonalization in each step of configuration recovery was performed with the SBD solver which can be obtained through the GitHub repository.65 The diagonalization in the SBD solver is parallelized with 48 CPUs used in the diagonalization of each subspace, where we use a total of 480 CPUs in 10 batches for each EWF cluster. The full sequence of configuration recovery, subspace construction, and diagonalization steps is summarized schematically in Fig. 8.
Subspace augmentation (ext-SQD): The extended SQD (ext-SQD) procedure was introduced to enable excited-state calculations based on SQD 5, 80 but may also be used to improve the accuracy of ground- and excited-state SQD calculations 6.
In ext-SQD we first take the electron configurations from the lowest energy batch on the last step of the configuration recovery procedure. We select the most dominant electron configurations based on the value of their configuration interaction coefficients (dominant if the threshold is above $1\cdot 10^{-5}$). Using the selected configurations, we then augment their subspaces by applying single excitations on each configuration within the selected set. The augmentation of the subspace was peformed with the PyCI software package.69 We use the augmented subspace to perform diagonalization within the given EWF cluster to obtain the ground state and energy of the Hamiltonian. Finally, we use the ground state produced after the diagonalization of the augmented subspace to calculate the reduced density matrices (RDM) of each EWF cluster.
@@ 5 Conclusions and Outlook
This study presents two important breakthroughs. First, the simulation of Trp-cage with STO-3G basis set is the largest example of embedding-based simulations within the configuration interaction paradigm and to our knowledge this is the largest case ever simulated exclusively within the configuration interaction framework. Trp-cage simulated with the STO-3G basis set results in simulation of 919 molecular orbitals (MOs), where the largest DMET and EWF simulations were reported only for systems explicitly including only up to 321 and 312 MOs, respectively.95, 56 Second, this represents the first demonstration of a system of this biochemical size being treated with quantum sampling on a quantum device as part of an embedding workflow, illustrating both the feasibility and the growing capability of hybrid quantum–classical electronic structure methods.
The present study paves the way for fragment-based solutions in quantum computing. Our approach is highly modular and allows for integration of novel quantum computing methodologies in the future. Hence, the present workflow will gradually mature in a synchronous manner with the maturation of quantum hardware and quantum algorithms. In the long-term, upon establishment of fault-tolerant quantum computing, the present fragment-based methodology can eventually adapt even more powerful fully-quantum methods such as quantum phase estimation (QPE) where perspectives of utilization of this method within embedding schemes were recently explored by Erakovic et al.17
Configuration interaction simulations accelerated with quantum computers can offer unique opportunities for simulations of complex chemical processes in proteins such as photochemical reactions and proton-coupled electron transfer45, 92. In the long-term, upon adaptation of more advanced future quantum hardware and quantum algorithms, this methodology potentially might offer a unified highly-accurate method for the treatment of any protein-based system beyond the scaling of classical computers.